1/4

Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников,

Дмитрий Николаевич Садовников,

Volga, Vol- ga (Stenka Rasin)

Stenka Rasin Sten’ka Razin Из-за острова на стрежень (Стенька Разин)

tradukita de N. N. 20 tradukita de N. N. 21 tradukita de Tr. Jacob Robbins

Bag om øen, på den

brede underskønne Vol-

gaflod Stenjka Rasins stolte

både ud på nye togter

stod.

In den Wellen hinter

Inseln Ziehen Kähne male-

risch, - |: Fangen leis an auf-

zuwachen, Bang ist jedes Ange-

sicht. :| 1. From beyond the woo-

ded island To the river wide and free Proudly sailed the arrow-

breasted Ships of Cossack yeoman-

ry.

1. Из-за острова на стрежень, На простор речной

волны, Выплывают

расписные, Острогрудые челны.

På den første

høvding Stenjka sidder selv i bådens

stavn, han på kamp og strid

ej tænker, har prinsessen i sin

favn.

Stenka Rasin vorn

als erster, Selig in der Trunken-

heit, |: Hält im Arme die

Prinzessin, Die er eben erst be-

freit. :| 2. On the first is Stenka

Razin With a princess at his si-

de, |: Drunken, holds a mar-

riage revel With his beautiful young

bride. :| 2. На переднем Стенька Разин, Обнявшись, сидит с княжной, Свадьбу новую справляет, Сам веселый и хмельной.

Men hans mænd er

Donkosakker, de er vrede, for de

ved, han har glemt dem

for en kvinde, for en nat i

kærlighed.

.................................................. .................................................. .................................................. ..................................................

3. But behind them rose a whisper,

“He has left his sword to

woo; |: One short night, and

Stenka Razin Has become a woman

too!” :| 3. А она, потупив

очи, Ни жива и ни мертва,

Молча слушает хмельные Атамановы слова.

Stenjka Rasin hører

latter, blodet koger i hans

bryst Over Volgas brede

strømme højt nu lyder Stenj-

kas røst.

.................................................. .................................................. .................................................. ..................................................

4. Stenka Razin hears the jeering

Of his discontented band, |: And the lovely Persian

princess He has circled with his

hand. :| 4. Позади их слышен

ропот:

- Нас на бабу променял! Только ночь с ней провозилс Сам наутро бабой

стал .

…  …  …  …

2/4

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

5. His black brows have come together

As the waves of anger rise, |: And the blood comes

rushing swiftly To his piercing, jet-black

eyes. :| 5. Этот ропот и насмешки Слышит грозный атаман, И могучею рукою Обнял персиянки

стан.

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

6. “I will give you all you ask for,

Life and heart, and head

and hand,” |: Echo rolls the pealing

thunder Of his voice across the

land. :| 6. Брови черные сошлися, Надвигается гроза. Буйной кровью налилис Атамановы глаза.

....................................... ....................................... ....................................... .......................................

Plötzlich tönt ein

dumpf Gemurre: Er verrät uns um ein

Weib, |: All der Seinen

Glück vergißt er Um geringen Zeit-

vertreib. :| 7. “Volga, Volga, mother

Volga, Deep and wide beneath

the sun, |: You have never seen a

present From the Cossack of the

Don. :|

7.  ”Ничего  не

пожалею, Буйну голову отдам!“ — Раздается голос

властный По окрестным берегам.

....................................... ....................................... ....................................... .......................................

Wolga, Wolga, Mut-

ter Wolga, Ruhig fließest du da-

hin, |: Ahnst nicht, was

ein Donkosake Fürchterliches hat

im Sinn. :| 8. And that peace might

rule as always All my free-born men and

brave, |: Volga, Volga, mother

Volga, Volga, make this girl a

grave.” :| 8. “Волга, Волга, мать родная, Волга, русская река, Не видала ты

подарка От донского казака!

”For at ingen splid

skal være mellem frie mænd,-

tag så Moder Volga, denne

kvinde!“ og hans arm om hen-

de lå.

.................................................. .................................................. .................................................. ..................................................

9. With a sudden, mighty movement,

Razin lifts the beauty

high, |: And he casts her where

the waters Of the Volga move and

sigh. :| 9. Чтобы не было раздора Между вольными людьми,

Волга, Волга, мать

родная, На, красавицу возьми!”

…  …  …  …

3/4

Og han ta’r med

stærke arme da sin brud og kaster

ned /: i de dybe, mørke

bølger, sine nætters salig-

hed. :/ Vorn als erster Stenka Rasin Hebt das Weib in wilder Wut, - |: Wolga, Wolga, nimm ein Opfer! Und er wirft sie in die Flut. :| 10. Now a silence like the

grave sinks To all those who stand to

see, |: And the battle-hardened

Cossacks Sink to weep on bended

knee. :| 10. Мощным взмахом поднимает Он красавицу

княжну И за борт ее бросает В набежавшую

волну.

“Men hvad fanden,

hvorfor hænger I med næbbet? Bliv

kun ved! Dans og syng de

røversange, syng for hendes

sjælefred!”

Und er sieht sie untergehen, Hört noch ihren Jammerschrei: |: Stirb als Opfer meiner Treue! Stenka Rasin, er ist

frei. :| 11. “Dance, you fool, and men, make merry! What has got into your

eyes? |: Let us thunder out a

chanty Of a place where beauty

lies.” :|

11.  “Что  ж

вы, братцы, приуныли? Эй, ты, Филька, черт, пляши! Грянем песню

удалую На помин ее души!..”

Frem fra øen - ud på

dybet stævner Stenjka Ra-

sin ud /: med den stærke,

stolte både, sejler fra sin døde

brud. :/ Und die Kähne ziehen weiter, Und die Kähne ziehen fort, |: Und die Wolga fließet weiter Über diesen Schreckensort.

12. From beyond the wooded island

To the river wide and free, |: Proudly sail the arrow-

breasted Ships of Cossack yeoman-

ry. :| 12. Из-за острова на стрежень, На простор речной

волны, Выплывают

расписные Острогрудые челны.

…  …  …  …

4/4

Traduko de la Ru-

sa  poemo  “Из-
за  острова  на

стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Danan de N. N. 20.

Arg-478-1037 (2010-03-08 12:16:20)

Aperinte en 1947 en la kant-libro “DLEA-Kantaro” eldonita de Dana Laborista Esperanto-Asocio (DLEA), n-ro 100 kaj en la kantlibro “Store Esperanto-Sangbog”, SEFO 1939, n-ro 31. Vidu la retejon esperantonia. dk sub http: // www. esperantonia. dk/ volga. htm .

Traduko de la Ru-

sa  poemo  “Из-
за  острова  на

стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Germanan de N. N. 21.

Arg-478-1038 (2010-03-09 15:31:43)

Mi, Manfred Retzlaff, trovis tiun ĉi germanigon de la kanto en http: // www. musicanet. org/ robokopp/ Lieder/ stenkara. html . Latinskriba aliskribo kaj laŭvorta germanigo troveblas en http: // www. kaikracht. de/ balalaika/ songs/ izza_bal. htm . Vidu ankaŭ la vikipedian retejon http: // de. wikipedia. org/ wiki/ Stenka_Rasin .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .

Verkinto de tiu ĉi Rusa poemo estas Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883).

Arg-478-927 (2010-03-07 18:59:55)

Pri la aŭtoro vidu ankaŭ la informojn sub http: // ru. wikipedia. org , la ruslingva parto de Wikipedia. Informoj pri la kanto, la melodio kaj diversaj versioj de la kanto estas troveblaj en la retejo http: // a-pesni. golosa. info/ popular20/ izzaostrova. htm .